OpenStudy (pottersheep):

Grade 12 Polynomial Functions help please ~

OpenStudy (pottersheep):

When 3x^2 + 10 x -3 is divided by x+ k , the remainder is 5. Determine values of k.

OpenStudy (anonymous):

according to the remainder theorem u can write\[3x^2+10x-3=(x+k)Q(x)+r\]

OpenStudy (anonymous):

u have r as 5

OpenStudy (anonymous):

so\[3x^2+10x-3=(x+k)Q(x)+5\]now let x=-k and solve for k

OpenStudy (anonymous):

\[3k^2-10k-3=(-k+k)Q(k)+5\]????

OpenStudy (pottersheep):

Thanks